OpenStudy (usukidoll):

Help! Calculus III problem Find the distance from the point to the plane (2,3,4), x+2y+2z =13

OpenStudy (goformit100):

Uae 3d - geometrical concept....

OpenStudy (usukidoll):

I haven't even learned that. The professor didn't explain it too good.

OpenStudy (experimentx):

you are doing multivariable calculus?

OpenStudy (usukidoll):

no. but it involves the usage of this formula.  where do I get the Point in this problem?

OpenStudy (usukidoll):

I don't know how to do that x(

OpenStudy (experimentx):

sorry ... passing through the point.

OpenStudy (experimentx):

take gradient of the plane ... that will give dc you normal.

OpenStudy (usukidoll):

gradient? is there a formula for that? or... is it based off of x+2y+2z=13?

OpenStudy (experimentx):

represent the line as (x-2)/1 = (y-3)/2 = (z-4)/3 = k

OpenStudy (usukidoll):

where did the 1, 2, and 3 come from?

OpenStudy (experimentx):

solve this equation with the plane x+2y+2z =13 to get x, y, z